OpenStudy (anonymous):

- Students in class are exchanging pencils. 1 student = 0 exchanges 2 students = 2 exchanges 3 students = 6 exchanges 4 students = 12 exchanges how many exhcanges for 7 sutudents? *I know the answer to be: 42 exchanges... but how to get it??? I knwo there is a formula, but I dont knwo how is it and how to use it..* help

Parth (parthkohli):

1 * 0 = 0 2 * 1 = 2 3 * 2 = 6 4 * 3 = 12 5 * 4 = 20 6 * 5 = 30 7 * 6 = 42

OpenStudy (anonymous):

yes, thats how i got it... but i know there is a formula... something like an=2n-1+(n-1) or something like that... (not sure if its that one.... and i know how to use that formula... nor i dont kknow what the formula menas.. :/

OpenStudy (opcode):

Okay. Math in the morning ^_^. There's a pattern. 1 student = 0 exchanges 2 students = 2 exchanges 3 students = 6 exchanges 4 students = 12 exchanges S students = X exchanges S students = X exchanges 7 students = 42 exchanges S = Students X = Exchanges The pattern is how many number of students there is minus one from it then times it by that number. So: S - 1 = Number to multiple by let's call it (N). S * N = X Let's test it: First say there are 10 students. 10 - 1 = 9 10 * 9 = 90 Exchanges. Get it?

OpenStudy (anonymous):

but why S-1???

Parth (parthkohli):

Nobody knows...

OpenStudy (opcode):

Because. The pattern shows that the number multiplied by the students is always one less than it.

OpenStudy (anonymous):

still dont get it.. (sorry for the persistance..) i wanna learn math like for real..

OpenStudy (opcode):

So the formula is: (S - 1) = N S * N = X S = Students X = Exchanges N = Number to multiple on "S".

OpenStudy (opcode):

Say there is: 120 students = ? exchanges Now use the formula: (S - 1) = N S = 120 So (120 - 1) = 119 N = 119 Now for step two: S * N = X We know what S and N are now so plug in the values: 120 * 119 = 14,280 14,280 = X And again: S = Students X = Exchanges N = Number to multiple on "S".

OpenStudy (anonymous):

yeah.. i get that... i just didnt know why N = S-1 ... but its ok tho... thanks man! :)

OpenStudy (anonymous):

u are a really good helper :- )

OpenStudy (opcode):

So the formula is: (S - 1) = N <-- S is for Students (Given) N is for the Number to multiple on "S". IN the pattern to get N. In this case we minus by "1". Then leading us to: S * N = X S = Students X = Exchanges N = Number to multiple on "S". And no problem ^_^.

Parth (parthkohli):

I am waiting to help like Opcode. :-(

OpenStudy (opcode):

Help me with what O.o? My stuff is just memorizing things like postulates and theorems. Memory with numbers = Head Ache. (For me at lest).

OpenStudy (anonymous):

ahah...